FOURIER TRANSFORM EMISSION SPECTROSCOPY OF THE B^{2}Σ^{+}–X^{2}Σ^{+} (VIOLET) SYSTEM OF 13C^{14}N

\underline{R. S. RAM} and P. F. BERNATH, Department of Chemistry, University of York, Heslington, York YO10 5DD; also at: Department of Chemistry, University of Arizona, Tucson, AZ 85721.

Emission spectra of the B^{2}Σ^{+}–X^{2}Σ^{+} transition of ^{13}C^{14}N have been observed at high resolution using the Fourier transform spectrometer associated with the McMath-Pierce Solar Telescope of the National Solar Observatory. The spectra have been measured in the 21000–30000 cm^{−1} region and a total of 52 vibrational bands involving vibrational levels up to v = 15 of the ground and excited states have been rotationally analyzed to provide a much improved set of spectroscopic constants. The results of the present analysis should prove useful in the identification of additional ^{13}C^{14}N lines in comets and cool stars, and will help in the determination of the ^{12}C/^{13}C abundance ratio. The observation of a number of highly-excited vibrational bands of the A^{2}Π–X^{2}Σ^{+} transition as well as a few bands of the B^{2}Σ^{+}–A^{2}Π transition will also be reported.